# Bifurcations and multistability in inducible three-gene toggle switch networks

### v1

@@ Acknowledgements.
Appendix A Analytic bifurcations for the two inducer model
In this appendix, we show analytically that the linear thresholds separating $(c_{1},c_{2})$ parameter space into tristable, bistable, and monostable regions in Section III.2 are the same as those derived for the single inducer model in Section III.1.
The first key bifurcation indicator with respect to $c_{1}$ is the loss of the $\bar{R}_{1}$-dominant stable state in transitioning from tristable to bistable dynamics. Assuming that $\bar{R}_{2,3}^{n}\rightarrow 0$ as in the single inducer setting, Eqns. 3133 indicate a steady state solution
@@ Appendix A Analytic bifurcations for the two inducer model
Since our assumption of the relative expression levels requires $\bar{R}_{2}=\bar{R}_{3}\leq\varepsilon$, it follows that
@@
and thus that the $\bar{R}_{1}$ dominant steady state exists for
@@
which is the same as the result derived in Eqn. 29 in the case of the single inducer.
We can derive the same result for $p_{\text{act}}(c_{2})$ by determining the bifurcation point with respect to $c_{2}$ at which the system no longer stabilizes to an $\bar{R}_{2}$-dominant state.
Appendix B Non-exclusive repressor-activator binding in the three-gene toggle switch
@@
Fig. 22 shows the thermodynamic states, weights and rates related to non-exclusive binding. Note that these include the same states (and corresponding rates) describing competitive binding in Fig. 12(A), but with three additional states in which an activator and either one or two repressors can be bound simultaneously at the relevant gene regulatory site. In these additional states, the presence of repressor(s) dampens expression from its activated rate to a+bda+b-d. The states and weights depicted translate into a corresponding system of equations for the evolution of expression. Omitting the role of effector, the change in $R_{1}$ expression, for instance, is
@@
with $\bar{d}\leq\bar{a}+\bar{b}$, $K^{(2)}=K_{2}/K_{1}$, and $K^{(3)}=K_{3}/K_{1}$. As in Section V.1.1, our analysis for this model assumes equal binding affinities such that $K^{(2)}=K^{(3)}=1$, and assumes a maximal production rate $\bar{a}+\bar{b}=2$.
@@
First, Fig. 22 defines effector activity as described by Case (A) of Fig. 11, where the active concentration of a protein indexed by $i$ is $p_{\rm act}(c)\bar{R}_{i}$ regardless of its function as a repressor or an activator. Therefore, for the system containing one effector at concentration $c$ that targets $\bar{R}_{1}$, the dimensionless Eqns. 5658 in induction Case (A) become
@@
Evaluating Eqns. 5961 for fixed points, we obtain the results discussed in Section V.2.1.
An alternative interpretation for effector activity, described by Case (B) of Fig. 11, transforms the available states and weights to those in Fig. 23, where the concentration of $\bar{R}_{i}$ is scaled by $p_{\rm act}(c)$ when repressing expression of other proteins, and scaled by $1-p_{\rm act}(c)$ when activating its own expression. For the system containing only one effector at concentration $c$ targetting $\bar{R}_{1}$ activity, Eqns. 5658 thus become
@@
Evaluating Eqns. 6264 for fixed points, we obtain the results discussed in Section V.2.2.

### v2

@@ Acknowledgements.
Appendix A Biologically permissible ranges for MWC parameters in allosteric induction
To determine the biologically permissible ranges for the MWC parameters defined in Eqn. 6 ($K_{I}$, $K_{A}$, and $\Delta\epsilon$), we assume that the absence of inducer does not imply the complete absence of inactive repressor, given that there may be some natural fractional error $q$. We suppose, then, that in the absence of inducer,
$p_{\text{act}}(c=0)\geq 1-q,$ (45)
or alternatively,
Rearranging Eqn. 46, we obtain a condition for the energy difference $\Delta\epsilon$:
$e^{-\beta\Delta\epsilon}\leq\frac{q}{1-q}.$ (47)
It follows that for the probability of activity to saturate in the absence of inducer ($q=0$), the energy of the inactive repressor state must far exceed that of the active state so that $\Delta\epsilon\rightarrow\infty$ and the repressor is always active.
Gene expression is generally “leaky,” meaning that there is typically a basal level of expression even in the presence of a large repressor concentration. We assume that this leakiness is also reflected by the fraction $q$ such that a fraction of repressors less than $q$ are active in the limit of infinite inducer concentration, with
$=\frac{(K_{I})^{m}}{(K_{I})^{m}+e^{-\beta\Delta\epsilon}(K_{A})^{m}}\leq q.$ (48)
Rearranging Eqn. 48, we obtain a second condition
Eqns. 47 and 49 establish bounds on the dissociation constants $K_{A,\>I}$ and the energy difference $\Delta\epsilon$ for a given “error” $q$ and number of binding sites $m$.
Suppose, for example, that less than $10\%$ ($q=0.1$) of repressors are active in the presence of an infinite inducer concentration. MWC parameters are then bounded by
$e^{-\beta\Delta\epsilon}\leq\frac{1}{9}$ (50)
and
$\Big(\frac{K_{A}}{K_{I}}\Big)^{m}\geq 9e^{\beta\Delta\epsilon}.$ (51)
We use these bounds to determine a permissible range of $K_{I}$ given fixed $K_{A}=150\>\mu M$ in Fig. 3(B).
Appendix B Analytic bifurcations for the two inducer model
In this appendix, we show analytically that the linear thresholds separating the $(c_{1},c_{2})$ parameter space into tristable, bistable, and monostable regions in Section III.2 are the same as those in Section III.1.
The key bifurcation with respect to $c_{1}$ is the loss of the $\bar{R}_{1}$-dominant stable state in transitioning from tristable to bistable dynamics. Assuming that $\bar{R}_{2,3}^{n}\rightarrow 0$ as in the single inducer setting, Eqns. 2426 indicate a steady state solution
@@ Appendix B Analytic bifurcations for the two inducer model
Since our assumptions require $\bar{R}_{2}=\bar{R}_{3}\leq\varepsilon$,
@@
and thus the $\bar{R}_{1}$ dominant steady state exists for
@@
the same as derived for a single inducer in Eqn. 22.
We derive the same bound for $p_{\text{act}}(c_{2})$ by determining the bifurcation point with respect to $c_{2}$ at which the system no longer stabilizes to an $\bar{R}_{2}$-dominant state.
Appendix C Non-exclusive repressor-activator binding in the three-gene toggle switch
@@
Fig. 22 shows the thermodynamic states, weights and rates related to non-exclusive binding. Note that these include the same states describing competitive binding in Fig. 12(A), but with three additional states in which an activator and either one or two repressors can be bound simultaneously at the relevant gene regulatory site. In these additional states, the presence of repressor(s) dampens expression to a+bda+b-d. The states and weights translate into a corresponding system of equations for the evolution of expression. Omitting the role of effector, the change in $R_{1}$ expression, for instance, is
@@
with $\bar{d}\leq\bar{a}+\bar{b}$, $K^{(2)}=K_{2}/K_{1}$, and $K^{(3)}=K_{3}/K_{1}$. As in Section V.1.1, our analysis in Section V.2 assumes equal binding affinities such that $K^{(2)}=K^{(3)}=1$, and a maximal production rate $\bar{a}+\bar{b}=2$.
@@
First, Fig. 22 defines effector activity as described by Case I of Fig. 11(A), where the active concentration of a protein indexed by $i$ is $p_{\rm act}(c)\bar{R}_{i}$ regardless of its function as a repressor or an activator. For the system containing one effector at concentration $c$ that targets $\bar{R}_{1}$, the dimensionless Eqns. 5658 in induction Case I become
@@
Evaluating the fixed points of Eqns. 5961, we obtain the results discussed in Section V.2.1.
An alternative interpretation for effector activity, described by Case II of Fig. 11(B), transforms the available states and weights to those in Fig. 23, where the concentration of $\bar{R}_{i}$ is scaled by $p_{\rm act}(c)$ when repressing expression of other proteins, and scaled by $1-p_{\rm act}(c)$ when activating its own expression. Eqns. 5658 then become
@@
$\frac{d\bar{R}_{3}}{d\bar{t}}$ $=\frac{\bar{a}}{[1+[p_{\rm act}(c)\bar{R}_{1}]^{n}][1+\bar{R}_{2}^{n}]}+\frac{\bar{b}\bar{R}_{3}^{n}}{1+\bar{R}_{3}^{n}}+\frac{(\bar{a}-\bar{d})\bar{R}_{3}^{n}\Big([p_{\rm act}(c)\bar{R}_{1}]^{n}+\bar{R}_{2}^{n}+[p_{\rm act}(c)\bar{R}_{1}]^{n}\bar{R}_{2}^{n}\Big)}{[1+[p_{\rm act}(c)\bar{R}_{1}]^{n}][1+\bar{R}_{2}^{n}][1+\bar{R}_{3}^{n}]}-\bar{R}_{3},$ (64)
describing a system containing only one effector at concentration $c$ targeting $\bar{R}_{1}$ activity. Evaluating the fixed points of Eqns. 6264, we obtain the results discussed in Section V.2.2.
Appendix D Code availability
All Matlab code used to generate graphs in figures throughout this work are available  [^1] Supporting Matlab code is available on GitHub at github.com/Rebecca-J-Rousseau/inducible3genetoggle.